# Aristotle’s Second-Order Logic

### v8

@@ 7. $\exists xPx\&Qx$ $\exists^{1}E$ 2-6
We propose that for Aristotle proof by ekthesis involved (essentially) the use of $\exists^{1}E$ in conjunction with other rules, perhaps mainly $\exists^{1}$.
@@
which is shown constructively by effectively producing such a $f$. So what is explicitly proven in Euclid is $\forall xT(x,fy)$, that for every segment $\overline{AB}$ the construction $f(\overline{AB})$ produces an equilateral triangle. It would seem that Proclus offers evidence that the ancients anticipated modern intuitionistic or constructivism interpretations of the existential quantifier666see[8] for further discussion on the constructivist interpretation of the existential quantification in Euclid. In appendix 3 we sketch how one might go about formalizing Euclid in ASOL in a way which seems natural and which also seems to fit within an intuitionistic or constructivist framework. in situations of multiple generality.
@@
Sentences of the form $\forall x\exists yRxy$ are found in Aristotle. The (modern) concept of function corresponds to the case $\forall x\exists^{1}yRxy$ and is found embedded in Aristotle’s Physics. It appears in context of proofs in expressions in the grammatical form ’the $X$ of $Y$’ (i.e. the time taken of a given motion) wherein it is understood that $\forall x\exists^{1}yRxy$ is assumed to hold. The proofs sometimes deploy $\exists^{1}$ on such expressions and this takes the form ’let $A$ be the $X$ of $Y$’ (for example: let $A$ be the time taken by body $X$ in motion $Y$). In the Physics there is an implied functional correspondence, for a given body, between all the continua of motion, space, time and magnitude. A division of one of these corresponds to a division of the others. Aristotle typically presents arguments involving a body $A$ moving in time $T$ together with the construction: ’let $S$ be the magnitude moved by $A$ in time $T$’, i.e. ’the $S$ of $A$ in $T$’ which supposes something like $\forall xt\exists^{1}yMov\,xyt$. The principle manner in which such suppositions are used corresponds to the deployment of $\exists^{1}$ together with other rules - which is our proposed interpretation of ekthesis. This technical term can be read as a sign that for Aristotle the deployment of the natural deduction rules for the existential quantifier was conscious and reflected.
@@
There are not many detailed mathematical proofs in the surviving Aristotelian corpus but the geometrically grounded proofs in the Physics and De Caelo et Mundi (see [6] for a good discussion of the mathematics of Aristotle) leave little doubt as to the presence of sentences with embedded quantifiers. For instance in Book VII of the Physics777241a25 in the proof we discussed that everything that moves is moved by something, hapan to kinoumenon hupo tinos anagkê kinesthai, Aristotle makes use of this axiom:
@@
We conclude this section by presenting some evidence for the use of quantifier rules over non-monadic predicates and situations embedded quantifiers in Sextus Empiricus. Note that we are evidently not implying that the mere presence of expressions involving multiple generality implies that the ancients had a mechanism for dealing with inferences over such kinds of expression. Rather we are presenting evidence from Sextus of inferences involving multiple generality presented explicitly as inferences888such inferences, if correct, are no small achievement if one bears in mind how often even the most basic fallacies in propositional logic are committed by people who we would not expect to exhibit faulty reasoning. Arguably natural language competency does necessarily imply logical competency. Showing that multiple generality could have been expressed among the Stoics by a systematic grammatical regimentation as argued in [5] is certainly an important step towards arguing for the historical presence of a Stoic quantifier logic. . It is likely that in late antiquity and the Hellenistic era logic become more syncretic and that there was substantial borrowing between the different schools. Thus even if the passage of Sextus in question seems to primarily concern Stoic logic it is probable that it also reflects indirectly later Peripatetic logic as well in particular with regards to the hypothetical syllogism expounded by Galen. Consider Sextus’ Outlines II, v, 23 (discussing a view of Democritus)[19]:
@@
We have the difference relation $\delta AB$ read as ’$A$ is a difference of genus $B$121212 A careful reading of the text (specially Book VI) reveals that Aristotle uses ’difference’ in three different ways: the difference of a species (relative to a genus or a definition), coordinate differences (of the same genus) and differences that ’belong’ to a given genus. We consider the third to be the most fundamental and define the two previous notions of difference in terms of this last one.. For property [^13] The topics of property set the stage for the topics of definition which is the culmination of the Topics. The notions deployed for property are basically the same one as those deployed for definition but is a more straightforward way. There is a much greater laxness in what can count as a property as compared to definition. We might say that property represents of a more archaic form of definition before the requirement of the a regimentation into a (proximate) genus and difference was adopted. The basic idea is that the property of a term $A$ cannot involve $A$ itself (non-circularity), cannot employ less known terms and cannot contain redundancies. $\iota AB$ read ’$B$ is a property of $A$’ we have
@@
It is important to note that equality here is not to be read syntactically but semantically. Thus $D=BC$ is to be read as ’the concept $D$ has as parts concepts $B$ and $C$’, the paradigm being ’rational animal’. We hold to be a fundamental problem whether a term can be said to be equal to its definition (with this notion of equality) and whether Leibniz’s law holds for this notion of equality. It could be that two distinct kinds of equality are required. It is true however that definition though the substitution of a term by its definition features in several topics171717152a-152b is very relevant for term equality. Aristotle announces topics which express the transitivity of $=$ as well as topics recalling Leibniz’s law. In 153a Aristotle seems to say that there is equality between a term and its definition but this is in itself not enough to show that it is a definition. The definiens must satisfy further conditions. The passage 140b32-141a1 is also interesting. Aristotle states that a definition is identical to the term it defines: esti de tauton tô(i) anthrôpô(i) zô(i)on pezon dipoun. However in 143a15-28 were Aristotle appears to admit that either giving the immediate genus and the difference or a non-immediate genus and all corresponding differences (the ’intension’) as both valid forms of definition which leads us to inquire exactly what the nature of the purported ’uniqueness’ of the definition. Does this means that both these forms are to be regarded as the same definition ?. From this last definition we can evidently derive (in ASOL) numerous refutation topics - and these are indeed topics in book VI. As Aristotle says: it is much easier to refute a definition than to establish one.
A standard view would be that if $hA(DC)$ then $C$ must be uniquely determined (this in fact follows from the above axioms). There is however a problem suggested by the passage 148b dealing with ’complex terms’181818this notion of ’complex term’ appears to be primarily grammatical and not logical or semantic as is our concatenation operation, however the precise relation between the two remains illusive as does the interpretation of the passage in question. Are we to understand the crucial non-circularity property of the definiens in semantic or syntactic terms ? But semantically does it even make sense to have a situation in which $A=AB$
@@
In the passages that follow Aristotle discusses the necessary agreement between the case governed for $\gamma_{2}$. For example if $\gamma_{2}RS$ then if $R$ takes the genitive then so does $S$. More interesting from a logical point of view is the passage mentioning the number of complements, this suggests that Aristotle was conscious of the necessity of the general agreement between the arity of the predicate terms in the species-genus relation, so that we should consider a whole sequence of such relations $\gamma_{1}=\gamma,\gamma_{2},...,\gamma_{i},....$ where $\gamma_{n}$ serves for $n$-ary predicate variables. Aristotle uses the example of the trinary relations ’donation’ and ’debt’. If we look at the early and middle Platonic dialogues we find that relation plays a much larger and sophisticated role than in does in Aristotle’s Organon. The logic of relations does not appear to be very worked at our and was perhaps consciously and deliberately neglected. Indeed there are some natural questions that could be asked (both now and in antiquity). For example: how do definitions work for relations ? What are the differences and how are they joined to the respective proximate genera ? The search for the definition of relatives clearly figures in the Platonic dialogues, for example in the Lysis the definition of friendship is sought after. A definition of a relation can necessarily involve sentences of arbitrary complexity involving other relations which do not fall under the general scheme for monadic predicate terms of book VI of the Topics (at least for its most simple familiar cases). Indeed traces of such unorthodox definitions are to be founded scattered throughout this work191919for example the definition of Justice as Courage and Temperance (150a-151b).. Even the most basic relations such as that of ’being a paternal grandfather’ do not seem to be organized neatly into a proximate genus $+$ difference format. In the Lysis the binary relations of friendship $\Phi$ and love $L$ are considered. The following hypothetical definition is put forward:
@@
In (1) and (2) the genera seem to be agôgê and summetria while eis ousian and thermôn kai psukhôn would correspond to prepositional phrase postmodifiers in English. In (6) it appears that the head or genus is hugron which is affected by two modifiers, the adjective apepton and the non-finite clause equivalent to a relative clause prôton apo trophês were the infinitive (probably ’generating’) is omitted. While a common noun + adjective definition for secondary substances can be construed as a conjunction of monadic predicate constants $G$ and $D$ yielding a definition term $\lambda x(Gx\&Dx)$ the situation is more delicate for other kinds of postmodifiers. One approach is to consider the case in which the genus or head is relation. For instance agôgê (coming to be, arising) is modified to ’coming to be according to essence’. This perhaps would be $\lambda Ax\top$ where $\top$ is the supreme genera for ousia.
@@
In (20) ’that which has three dimensions’ we have an example of a noun-phrase having as head a demonstrative pronoun with a relative clause postmodifier. The passage 140b32-141a1 is interesting. Aristotle’s point here seems to be that postmodifiers are not predicative so that a term may be predicated of a noun phrase (kata tou sumpantos) with a modifier employing the same term without redundancy (hapaks hê kategoria ginetai), for example in :a biped animal is biped. Finally we whould not fail to mention that ASOL is adequate to treat propositional attitudes through the use of second-order predicates involving 00-ary predicate terms (which correspond to propositions) and the $lambda$-construction for an empty list of variables as standard in SOL. For example $a$ knows that $Ab$ could be expressed as $Ka(\lambda(Ab))$. It would be interesting to look for instances of propositional attitudes in the Organon.

### v11

@@ 7. $\exists xPx\&Qx$ $\exists^{1}E$ 2-6
We propose that for Aristotle proof by ekthesis involved (essentially) the use of $\exists^{1}E$ in conjunction with other rules, perhaps mainly $\exists^{1}$. However a case can be made that Aristotle deployed reasoning in cases in which $\exists^{I}$ would be deployed which amounts to considering Skolemizations or Choice Functions of the corresponding formulas.
@@
which is shown constructively by effectively producing such a $f$. So what is explicitly proven in Euclid is $\forall xT(x,fy)$, that for every segment $\overline{AB}$ the construction $f(\overline{AB})$ produces an equilateral triangle. It would seem that Proclus offers evidence that the ancients anticipated modern intuitionistic or constructivism interpretations of the existential quantifier666see[11] for further discussion on the constructivist interpretation of the existential quantification in Euclid. In appendix 3 we sketch how one might go about formalizing Euclid in ASOL in a way which seems natural and which also seems to fit within an intuitionistic or constructivist framework. in situations of multiple generality. It could be that Skolemization or Choice Functions might be quite an important aspect of the logical treatment of propositions of the form $\forall X\exists Y\phi(X,Y)$ (here $X$ and $Y$ are either individual variables or predicate term variables) in Aristotle and other ancient authors, as an alternative to the deployment of natural deduction rule $\exists^{1}E$. For $\forall X\exists Y\phi(X,Y)$ we would have a choice function $j(X)$, where $\phi(X,j(X))$. This is like an indefinite dependent version $\epsilon$ of the Peano operator and $j(X)$ can be read as a $Y$ such that $\phi(X,Y)$. We can write $j(X)=\epsilon Y\phi(X,Y)$. In this way we interpret this constructions: let $a$ be a $Y$ such that $\phi(X,Y)$ or in more natural language form: let $a$ be a $\phi$ of $X$.
And we can consider this to include the case of $\exists X\phi(X)$ where we deploy the indefinite version of the Peano operator (which corresponds to a definite article): $a=\epsilon X\phi(X)$, let $a$ be a $\phi$.
Consider proposition 2777Given two unequal magnitudes, it is possible to find two unequal straight lines such that the greater straight line has to the less a ratio less than the greater magnitude has to the less (Heath tr.). of Archimedes’s On the Sphere and Cylinder (Book I). Archimedes assumes that given quantities $A<B$ there is natural number $n$ such that $n\bullet A>B$ where $n\bullet A$ means the quantity obtained by adding $A$ $n$ times to itself. Thus this axiom reads $\forall A,B\exists n(A<B\rightarrow(n\bullet A>B))$. The way Archimedes uses the axiom in the proof [^8] The relevant part of the proof reads (Heath tr.): Then, if $CA$ is added to itself a sufficient number of times, the sum will exceed $D$. Let $AF$ be this sum. This is equivalent to: let $AF$ be a multiple of $CA$ larger than $D$. is far more suggestive of the deployment of a Choice Function $j$ for the axiom than $\exists^{1}E$. Archimedes takes directly a $C=j(A,B,n)\bullet A$ where $j$ is a Choice Function for the axiom.
We could read the ekthesis in 25a14-18 examined above as follows (and similarly for 28a24-27):
1. $\forall x(Bx\rightarrow\sim Ax)$ Hyp
2. $\sim(\forall x(Ax\rightarrow\sim Bx))$ Hyp
3. $\exists xAx\&Bx$ 2
4. $c=\epsilon x(Ax\&Bx)$ 3
5. $Ac\&Bc$ by the expected axiom for $\epsilon$ and 3
6. $\bot$ 1,5
7. $\forall x(Ax\rightarrow\sim Bx)$ $\bot_{c}$ 2, 6
[…]
@@ 7. $\forall x(Ax\rightarrow\sim Bx)$ $\bot_{c}$ 2, 6
Sentences of the form $\forall x\exists yRxy$ abound in Aristotle. The (modern) concept of function corresponds to the case $\forall x\exists^{1}yRxy$ and is found embedded in Aristotle’s Physics. It appears in context of proofs in expressions in the grammatical form ’let $a$ be the $y$ such that $Rxy$’ (i.e. the time taken of a given motion) wherein it is understood that $\forall x\exists^{1}yRxy$ is assumed to hold. More generally we meet expressions of the form $\forall x,y\exists^{1}zMxyz$ and for a given $b$ and $m$ we find: let $t$ be the $z$ such that $Mbmz$ (which corresponds to the $\epsilon$ construction discussed above). For instance: let $t$ be the time taken by body $b$ in motion $m$. The proofs can be interpreted as deploying $\exists^{1}$, for instance using $\exists yRxy$ and hypothesis $Rxa$. In the Physics there is an implied functional correspondence, for a given body, between all the continua of motion, space, time and magnitude. A division of one of these corresponds to a division of the others. The principle manner in which such suppositions are used would seem to correspond to the deployment of $\exists^{1}$ together with other rules - which is our proposed interpretation of ekthesis. We propose that ekthesis may also be interpreted in terms of the $\epsilon$ operator (an interpretation we will develop in a future work). This technical term can be read as a sign that for Aristotle the deployment of the natural deduction rules for the existential quantifier was conscious and reflected. In this ’functional’ case of unique existential quantification we have unique choice, reflected in the corresponding definite article in the grammatical construction: let $y$ be the $x$ such that $Rxy$. In the case of $\exists^{1}X\phi(X)$ we have the usual Peano operator $\iota X\phi(X)$. For the situation $\forall X\exists^{1}Y\phi(X,Y)$ we can denote this by $j(X)=\iota Y\phi(X,Y)$. This corresponds to a genitive: $j(X)$ is read as $X$’s $\phi$ while $j(X)=\epsilon Y\phi(X,Y)$ is read as a $\phi$ of $X$999to simplify the discussion we have overlooked the fact that in Aristotle instead of sentences of the form $\exists X\phi(X)$ and $\forall X\exists Y\phi(X,Y)$ we mainly meet restricted versions such as $\exists X(\psi(X)\&\phi(X))$ and $\forall X\exists Y(\psi(X)\rightarrow(\phi(X,Y))$. For example: ’there is man who knows grammar’ and ’every man has a friend’. The $\epsilon$ construction would then correspond to: ’let $m$ be a man who knows grammar’ and ’let $f$ be a friend of $m$’. See also the discussion in [7] regarding the tis + ekeinos construction for such cases of multiple generality. .
@@
There are not many detailed mathematical proofs in the surviving Aristotelian corpus but the geometrically grounded proofs in the Physics and De Caelo et Mundi (see [9] for an overview of the mathematics of Aristotle) leave little doubt as to the presence of sentences with embedded quantifiers. For instance in Book VII of the Physics101010241a25 in the proof we discussed that everything that moves is moved by something, hapan to kinoumenon hupo tinos anagkê kinesthai, Aristotle makes use of this axiom:
@@
But we could also reconstruct the argument using the $\epsilon$ interpretation of ekthesis as seen above. In fact since the axiom can be strengthened to every man has a unique father the argument can be expressed very simply using the generalization of the Peano operator considered previously.
We conclude this section by presenting some evidence for the use of quantifier rules over non-monadic predicates and situations embedded quantifiers in Sextus Empiricus. Note that we are evidently not implying that the mere presence of expressions involving multiple generality implies that the ancients had a mechanism for dealing with inferences over such kinds of expression. Rather we are presenting evidence from Sextus of inferences involving multiple generality presented explicitly as inferences111111such inferences, if correct, are no small achievement if one bears in mind how often even the most basic fallacies in propositional logic are committed by people who we would not expect to exhibit faulty reasoning. Arguably natural language competency does necessarily imply logical competency. Showing that multiple generality could have been expressed among the Stoics by a systematic grammatical regimentation as argued in [7] is certainly an important step towards arguing for the historical presence of a Stoic quantifier logic. . It is likely that in late antiquity and the Hellenistic era logic become more syncretic and that there was substantial borrowing between the different schools. Thus even if the passage of Sextus in question seems to primarily concern Stoic logic it is probable that it also reflects indirectly later Peripatetic logic as well in particular with regards to the hypothetical syllogism expounded by Galen. Consider Sextus’ Outlines II, v, 23 (discussing a view of Democritus)[22]:
@@
We have the difference relation $\delta AB$ read as ’$A$ is a difference of genus $B$151515 A careful reading of the text (specially Book VI) reveals that Aristotle uses ’difference’ in three different ways: the difference of a species (relative to a genus or a definition), coordinate differences (of the same genus) and differences that ’belong’ to a given genus. We consider the third to be the most fundamental and define the two previous notions of difference in terms of this last one.. For property [^16] The topics of property set the stage for the topics of definition which is the culmination of the Topics. The notions deployed for property are basically the same one as those deployed for definition but is a more straightforward way. There is a much greater laxness in what can count as a property as compared to definition. We might say that property represents of a more archaic form of definition before the requirement of the a regimentation into a (proximate) genus and difference was adopted. The basic idea is that the property of a term $A$ cannot involve $A$ itself (non-circularity), cannot employ less known terms and cannot contain redundancies. $\iota AB$ read ’$B$ is a property of $A$’ we have
@@
It is important to note that equality here is not to be read syntactically but semantically. Thus $D=BC$ is to be read as ’the concept $D$ has as parts concepts $B$ and $C$’, the paradigm being ’rational animal’. We hold to be a fundamental problem whether a term can be said to be equal to its definition (with this notion of equality) and whether Leibniz’s law holds for this notion of equality. It could be that two distinct kinds of equality are required. It is true however that definition though the substitution of a term by its definition features in several topics202020152a-152b is very relevant for term equality. Aristotle announces topics which express the transitivity of $=$ as well as topics recalling Leibniz’s law. In 153a Aristotle seems to say that there is equality between a term and its definition but this is in itself not enough to show that it is a definition. The definiens must satisfy further conditions. The passage 140b32-141a1 is also interesting. Aristotle states that a definition is identical to the term it defines: esti de tauton tô(i) anthrôpô(i) zô(i)on pezon dipoun. However in 143a15-28 where Aristotle appears to admit that either giving the immediate genus and the difference or a non-immediate genus and all corresponding differences (the ’intension’) as both valid forms of definition which leads us to inquire exactly what the nature of the purported ’uniqueness’ of the definition. Does this means that both these forms are to be regarded as the same definition ?. From this last definition we can evidently derive (in ASOL) numerous refutation topics - and these are indeed topics in book VI. As Aristotle says: it is much easier to refute a definition than to establish one.
A standard view would be that if $hA(DC)$ then $C$ must be uniquely determined (this in fact follows from the above axioms). There is however a problem suggested by the passage 148b dealing with ’complex terms’212121this notion of ’complex term’ appears to be primarily grammatical and not logical or semantic as is our concatenation operation, however the precise relation between the two remains illusive as does the interpretation of the passage in question. Are we to understand the crucial non-circularity property of the definiens in semantic or syntactic terms ? But semantically does it even make sense to have a situation in which $A=AB$
@@
It would be interesting to investigate the connection between the theory of relations in the Topics and the interpretation of ekthesis in terms of $\epsilon$ and the generalization of $\iota$ discussed in section 3.
In the passages that follow Aristotle discusses the necessary agreement between the case governed for $\gamma_{2}$. For example if $\gamma_{2}RS$ then if $R$ takes the genitive then so does $S$. More interesting from a logical point of view is the passage mentioning the number of complements, this suggests that Aristotle was conscious of the necessity of the general agreement between the arity of the predicate terms in the species-genus relation, so that we should consider a whole sequence of such relations $\gamma_{1}=\gamma,\gamma_{2},...,\gamma_{i},....$ where $\gamma_{n}$ serves for $n$-ary predicate variables. Aristotle uses the example of the trinary relations ’donation’ and ’debt’. If we look at the early and middle Platonic dialogues we find that relation plays a much larger and sophisticated role than in does in Aristotle’s Organon. The logic of relations does not appear to be very worked at our and was perhaps consciously and deliberately neglected. Indeed there are some natural questions that could be asked (both now and in antiquity). For example: how do definitions work for relations ? What are the differences and how are they joined to the respective proximate genera ? The search for the definition of relatives clearly figures in the Platonic dialogues, for example in the Lysis the definition of friendship is sought after. A definition of a relation can necessarily involve sentences of arbitrary complexity involving other relations which do not fall under the general scheme for monadic predicate terms of book VI of the Topics (at least for its most simple familiar cases). Indeed traces of such unorthodox definitions are to be founded scattered throughout this work222222for example the definition of Justice as Courage and Temperance (150a-151b).. Even the most basic relations such as that of ’being a paternal grandfather’ do not seem to be organized neatly into a proximate genus $+$ difference format. In the Lysis the binary relations of friendship $\Phi$ and love $L$ are considered. The following hypothetical definition is put forward:
@@
The above list provides conclusive evidence that a great variety of noun phrase constructions (and in particular with regards to the structure of the propositional phrases appearing as postmodifiers) were allowed by Aristotle. In (1) and (2) the genera seem to be agôgê and summetria while eis ousian and thermôn kai psukhôn would correspond to prepositional phrase postmodifiers in English. In (6) it appears that the head or genus is hugron which is affected by two modifiers, the adjective apepton and the non-finite clause equivalent to a relative clause prôton apo trophês were the infinitive (probably ’generating’) is omitted. While a common noun + adjective definition for secondary substances can be construed as a conjunction of monadic predicate constants $G$ and $D$ yielding a definition term $\lambda x(Gx\&Dx)$ the situation is more delicate for other kinds of postmodifiers. One approach is to consider the case in which the genus or head is relation. For instance agôgê (coming to be, arising) is modified to ’coming to be according to essence’. This perhaps would be $\lambda Ax\top$ where $\top$ is the supreme genera for ousia.
@@
In (20) ’that which has three dimensions’ we have an example of a noun-phrase having as head a demonstrative pronoun with a relative clause postmodifier. The passage 140b32-141a1 is interesting. Aristotle’s point here seems to be that postmodifiers are not predicative so that a term may be predicated of a noun phrase (kata tou sumpantos) with a modifier employing the same term without redundancy (hapaks hê kategoria ginetai), for example in :a biped animal is biped [^24] This might be interpreted as Aristotle pointing out the logical-grammatical difference between predication of a concept to a subject and the concept used to specify the subject (typically relative clause modifiers) as in ’the prime numbers’ vs. ’quantities which are numbers’. This distinction is parallel to the Fregean distinction between proposition and judgment. . Finally we whould not fail to mention that ASOL is adequate to treat propositional attitudes through the use of second-order predicates involving 00-ary predicate terms (which correspond to propositions) and the $\lambda$-construction for an empty list of variables as standard in SOL. For example $a$ knows that $Ab$ could be expressed as $Ka(\lambda(Ab))$. It would be interesting to look for instances of propositional attitudes in the Organon.